Kaitlyn is flipping a coin two times, and records the result of each flip. Which of the sets below is the entire sample space?

{HH,TT,HT,TH}


What is the probability of obtaining heads twice?


  


What is the probability of obtaining heads exactly once?


  


What is the probability of obtaining heads at least once?


  


What is the probability of obtaining no heads?


1
Expert's answer
2022-05-26T14:34:38-0400

The sample space of two tossed coins is:

S: {(H,H),(H,T,),(T,H),(T,T)}

Probability=(No. Of favourable outcomes)/(Total no. Of outcomes)

1-- Here, from the sample space,

P(getting 2 heads when a coin is tossed twice)= 1/4

2-- the probability that we get exactly one head is the probability that we get anyone of the following 2 outcomes: HT, TH. Each has probability 1/4, so the probability to get exactly one head in 2 flips is 1/4 + 1/4 = 1/2.

3-- In three of the four outcomes, a Heads appears, so probability of at least one head is 3/4.

4-- the probability of obtaining no heads (TT) out of 4 outcomes is 1/4.